# Identifying Constitutive Parameters for Complex Hyperelastic Materials using Physics-Informed Neural Networks

### v3

@@ 3.2 Identify constitutive parameters for the phototype problem
In the last subsection, we demonstrated the proposed PINNs framework can efficiently model the forward problem. In the following subsections, we apply the framework to identify the unknown constitutive parameters, $\bm{\theta}_{\text{const}}=(\mu,\lambda_{m}$), based on synthetic experimental data. Unlike conventional approaches for inverse problems Yan and Diebold (2021a, b), which formulate as an initial value or boundary value problem, PINNs take a distinctive route by minimizing the loss functions inherent in the constitutive relations. In this subsection, we first employed the PINNs framework to the prototype problem described in Section 2.2 (Fig.2(a)). We randomly selected 400 sampling points within the subdomain, $0.80.2<|X_{1}|<$ 0.8, to calculate the total loading force at each time step using Eq.(9). We tuned the weights of each loss to maximize the prediction performance with emphasis on $\mathcal{L}^{\text{inc}}$ and $\mathcal{L}^{\text{Loading}}$.
@@
Figure 3 illustrates the predicted values for $\mu$ and $\lambda_{m}$ for three sets of true values. In the training process, the displacements of 200 randomly selected speckles over 21 time steps serve as additional constraints, except for the green curves. Starting with initial values of $(\mu,\lambda_{m})_{\text{initial}}=(4.0,7.0)$, the model with the DIC data demonstrates better convergence in less than $10^{6}$ epochs, compared with the case without the DIC data (green curves). The convergence speed for the inverse problem is reasonable since the forward model also requires at least 10k training epochs to obtain a reasonably accurate deformation field prediction (Fig.2(b)). Interestingly, we also observed that the PINNs model had quicker convergence when predicting $\mu$ over $\lambda_{m}$, which first optimized the material parameters akin to a NH model. Subsequently, the model progressively learned the strain-hardening behaviors by converging $\lambda_{m}$ to the ground truth. This dynamic evolution in parameter optimization highlights the model’s capability to adapt and capture intricate material behaviors as training progresses. In this section, we investigated a relatively simple scenario involving a circular hole in the middle of the sample. In such case, we could derive the theoretical approximations by fitting the loading curving to the 1D uniaxial tension solutions, which yield reasonable predictions with $\mu^{est}\approx 0.94$ and $\lambda_{m}^{est}\approx 2.97$ when the true values are 1.0 and 3.0, respectively. However, most biological soft materials have much more intricate sample geometries, where theoretical estimations or other inverse techniques cannot readily fit their material parameters. In the subsequent section, we will apply the PINNs framework to a more complex scenario, demonstrating that the loading history data cannot sufficiently identify the parameters. Therefore, integrating DIC data is necessary for parameter inversion with the desired accuracy in handling intricate cases.
@@ 3.3 Identify constitutive parameters for complex geometries
The PDEs were evaluated within the region delineated by light grey shading, as depicted in Fig.4(a). Furthermore, we incorporated the displacement history of 200 randomly distributed speckles across 21 time steps. Due to the unknown material properties of the four circular inhomogeneities, we utilized the displacement field data obtained from synthetic DIC as BCs for these inhomogeneities. We tuned the weights of loss functions with emphasis on both speckle displacement and the loading history. These optimized weights account for the varying degrees of influence each constraint exerts on the optimization process, effectively guiding the model to converge toward accurate parameter values despite the challenges posed by the complex configuration. We have tested our model with a variety of initial values for the constitutive parameters in a reasonable range and all of them show similar convergence behavior. For illustration, we present the results with the initial values of $(\mu,\lambda_{m})_{\text{initial}}=(4.0,7.0)$ in Figs.4(b) and (c). As shown in Fig.4(b), the losses experience rapid reduction within the first 10k training epochs, followed by a more gradual decline to convergence. In Fig.4(c), we plotted the displacement outlines predicted by the PINNs framework and the corresponding FEM simulations. Remarkably, despite not explicitly modeling the inherent inhomogeneity, the PINNs adeptly predicted the deformation field based on synthetic DIC data. The predictions for $\mu$ and $\lambda_{m}$ are presented in Figs.4(d)&(e). We tested multiple initial values, all of which demonstrated similar convergence performance. Notably, there is a pronounced increase in $\lambda_{m}$ during the initial 180k training epochs, signifying convergence toward the NH limit. As the PDE loss and BC loss gradually diminished, the PINNs framework started to capture the strain-stiffening behavior.
@@ 4 Conclusions
We first verified our framework with the forward modeling on a phototype problem and found excellent agreement with FEM results. Then, we applied our framework for parameter inversion with the phototype problem and beyond. Our framework introduced in this paper is particularly suited for cases involving intricate sample geometries that cannot be readily addressed by theoretical estimations or other inverse techniques. Our model pinpointed the necessary parameters while ensuring errors were below 5%. It is particularly noteworthy that this impressive accuracy was maintained even when the framework was trained with an experimental noise level of 5% with sufficient full-field measurement density.
The current PINNs framework can be readily extended to other classes of hyperelastic models, such as Yeoh model Yeoh (1993), HGO model Holzapfel et al. (2000), Gent model Gent (1996), etc. For these constitutive relations emphasizing the strain-stiffening anisotropic behaviors, we anticipate they will show similar behavior to the AB model in that the PINNs converge to an NH limit in the first few epochs and gradually learn the stress-strain curves. In addition to the identification of the materials parameters for constitutive models, the current PINNs framework integrating with DIC data and loading history curves can be applied to extract other material properties, such as cohesive zone parameters Jin (2022); Wei et al. (2022), residual stress in thin films Rao et al. (2020); Chason and Guduru (2016), and the flow resistance Song et al. (2018).
@@
Funding
This work does not receive any external funding.

### v4

@@ 3.2 Identify constitutive parameters for the prototype problem
In the last subsection, we demonstrated the proposed PINNs framework can efficiently model the forward problem. In the following subsections, we apply the framework to identify the unknown constitutive parameters, $\bm{\theta}_{\text{const}}=(\mu,\lambda_{m}$), based on synthetic experimental data. Unlike conventional approaches for inverse problems [54, 55, 56, 57, 58], which formulate as an initial value or boundary value problem, PINNs take a distinctive route by minimizing the loss functions inherent in the constitutive relations. In this subsection, we first employed the PINNs framework to the prototype problem described in Section 2.2 (Fig.2(a)). We randomly selected 400 sampling points within the subdomain, $0.80.2<|X_{1}|<$ 0.8, to calculate the total loading force at each time step using Eq.(9). We tuned the weights of each loss to maximize the prediction performance with emphasis on $\mathcal{L}^{\text{inc}}$ and $\mathcal{L}^{\text{Loading}}$.
@@
Fig.3 illustrates the predicted values for $\mu$ and $\lambda_{m}$ for three sets of true values. In the training process, the displacements of 200 randomly selected speckles over 21 time steps serve as additional constraints, except for the green curves. Starting with initial values of $(\mu,\lambda_{m})_{\text{initial}}=(4.0,7.0)$, the model with the DIC data demonstrates better convergence in less than $10^{6}$ epochs, compared with the case without the DIC data (green curves). The convergence speed for the inverse problem is reasonable since the forward model also requires at least 10k training epochs to obtain a reasonably accurate deformation field prediction (Fig.2(b)). Interestingly, we also observed that the PINNs model had quicker convergence when predicting $\mu$ over $\lambda_{m}$, which first optimized the material parameters akin to a NH model. Subsequently, the model progressively learned the strain-hardening behaviors by converging $\lambda_{m}$ to the ground truth. This dynamic evolution in parameter optimization highlights the model’s capability to adapt and capture intricate material behaviors as training progresses. In this section, we investigated a relatively simple scenario involving a circular hole in the middle of the sample. In such case, we could derive the theoretical approximations by fitting the loading curving to the 1D uniaxial tension solutions, which yield reasonable predictions with $\mu^{est}\approx 0.94$ and $\lambda_{m}^{est}\approx 2.97$ when the true values are 1.0 and 3.0, respectively. However, most biological soft materials have much more intricate sample geometries, where theoretical estimations or other inverse techniques cannot readily fit their material parameters. In the subsequent section, we will apply the PINNs framework to a more complex scenario, demonstrating that the loading history data cannot sufficiently identify the parameters. Therefore, integrating DIC data is necessary for parameter inversion with the desired accuracy in handling intricate cases.
@@ 3.3 Identify constitutive parameters for complex geometries
The PDEs were evaluated within the region delineated by light grey shading, as depicted in Fig.4(a). Furthermore, we incorporated the displacement history of 200 randomly distributed speckles across 21 time steps. Due to the unknown material properties of the four circular inhomogeneities, we utilized the displacement field data obtained from synthetic DIC as BCs for these inhomogeneities. We tuned the weights of loss functions with emphasis on both speckle displacement and the loading history. These optimized weights account for the varying degrees of influence each constraint exerts on the optimization process, effectively guiding the model to converge toward accurate parameter values despite the challenges posed by the complex configuration. We have tested our model with a variety of initial values for the constitutive parameters in a reasonable range and all of them show similar convergence behavior. For illustration, we present the results with the initial values of $(\mu,\lambda_{m})_{\text{initial}}=(4.0,7.0)$ in Figs.4(b) and (c). As shown in Fig.4(b), the losses experience rapid reduction within the first 10k training epochs, followed by a more gradual decline to convergence. In Fig.4(c), we plotted the displacement outlines predicted by the PINNs framework and the corresponding FEM simulations. Remarkably, despite not explicitly modeling the inherent inhomogeneity, the PINNs adeptly predicted the deformation field based on synthetic DIC data. The predictions for $\mu$ and $\lambda_{m}$ are presented in Figs.4(d)&(e). We tested multiple initial values, all of which demonstrated similar convergence performance. Notably, there is a pronounced increase in $\lambda_{m}$ during the initial 180k training epochs, signifying convergence toward the NH limit. As the PDE loss and BC loss gradually diminished, the PINNs framework started to capture the strain-stiffening behavior. The convergence of our model remains stable across various initial values. However, the disparity between initial guesses and the ground truth impacts the rate of convergence. As depicted in Fig.3(b) and Fig.4(e), the smaller initial value of the $\lambda_{m}$ typically leads to faster convergence. This trend arises because the optimization process tends to overestimate $\lambda_{m}$ in order to minimize errors in the power term, and gradually adjusts its value towards the ground truth. Conversely, if the initial value of $\lambda_{m}$ is set too large, the overestimation becomes more pronounced, resulting in a longer time for convergence.
@@ 4 Conclusions
We first verified our framework with the forward modeling on a prototype problem and found excellent agreement with FEM results. Then, we applied our framework for parameter inversion with the prototype problem and beyond. Our framework introduced in this paper is particularly suited for cases involving intricate sample geometries that cannot be readily addressed by theoretical estimations or other inverse techniques. Our model pinpointed the necessary parameters while ensuring errors were below 5%. It is particularly noteworthy that this impressive accuracy was maintained even when the framework was trained with an experimental noise level of 5% with sufficient full-field measurement density.
The current PINNs framework can be readily extended to other classes of hyperelastic models, such as Yeoh model [59], HGO model [5], Gent model [60], etc. For these constitutive relations emphasizing the strain-stiffening anisotropic behaviors, we anticipate they will show similar behavior to the AB model in that the PINNs converge to an NH limit in the first few epochs and gradually learn the stress-strain curves. Furthermore, the present PINNs framework is versatile and can be extended to accommodate rate-dependent hyperelastic models, including the Bergstrom-Boyce model [61]. This adaptability is facilitated by the ability to reformulate the PDEs into an incremental form, similar to approaches used in the study of finite deformation plasticity [32]. By incorporating speckle information into the incremental formulation, we anticipate improvements in convergence speed and applicability to more complex sample structures. In addition to the identification of the materials parameters for constitutive models, the current PINNs framework integrating with DIC data and loading history curves can be applied to extract other material properties, such as cohesive zone parameters [17, 62], residual stress in thin films [63, 64], and the flow resistance [65].
Appendix: Identifying material parameters for samples with square inhomogeneities
The proposed PINNs model is versatile for identifying material parameters from samples with diverse geometries. As an example, we examined a case involving square inhomogeneities which are represented as stiff bodies, each with an edge length of 0.2 (Fig.A1). 200 random speckle points were added as BCs with no noise. As shown in Fig.A1, after 80k training epochs, the total loss is less than $10^{-4}$. The prediction has a relative error of $\frac{|\mu-\mu^{\text{True}}|}{\mu^{\text{True}}}=13.79\pm 0.66\%$, $\frac{|\lambda_{m}-\lambda_{m}^{\text{True}}|}{\lambda_{m}^{\text{True}}}=2.06\pm 1.69\%$. Compared to the samples with circular holes (Fig.4), the stress field near the corners of the square inhomogeneities exhibits singularity, resulting in slower convergence and lower prediction accuracy especially for $\mu$. Despite the complexity of these defects, the current PINNs framework remains stable, providing reasonable predictions. To enhance the accuracy due to the stress concentration, a multiscale approach which can accurately represent the local stress field can be employed in the future work [53].
Figure A1: Constitutive parameter identification for a rectangular sample with four square inhomogeneities. (a) Schematic of the sample’s geometry and dimension: the light grey region represents AB material with the material parameters $(\mu,\lambda_{\text{m}})$, while the dark grey region represents incompressible NH material. (b) Loss as a function of training epochs. (c) The prediction of $\mu$ as a function of the training epochs. (d) The prediction of $\lambda_{\text{m}}$ as a function of the training epochs. The dashed line is the predictions from PINNs and the solid line is the ground truth.